# Deep End-to-end Adaptive k-Space Sampling, Reconstruction, and Registration for Dynamic MRI

### v1

@@ 0.C.1 Additional Tables
$\alpha$ $\beta$ SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$)
vSHARP Adaptive Learned 1 1 0.877 $\pm$ 0.050 31.68 $\pm$ 2.86 0.037 $\pm$ 0.019 0.867 $\pm$ 0.052 31.41 $\pm$ 2.76 0.038 $\pm$ 0.019 0.854 $\pm$ 0.054 31.01 $\pm$ 2.69 0.042 $\pm$ 0.020
vSHARP Adaptive Learned 1 1 0.880 $\pm$ 0.048 31.95 $\pm$ 2.81 0.034 $\pm$ 0.018 0.872 $\pm$ 0.049 31.75 $\pm$ 2.70 0.035 $\pm$ 0.017 0.862 $\pm$ 0.052 31.41 $\pm$ 2.64 0.038 $\pm$ 0.018
vSHARP Adaptive Learned 1 1 0.882 $\pm$ 0.049 31.90 $\pm$ 2.92 0.035 $\pm$ 0.019 0.874 $\pm$ 0.051 31.75 $\pm$ 2.84 0.036 $\pm$ 0.019 0.862 $\pm$ 0.054 31.40 $\pm$ 2.75 0.039 $\pm$ 0.020
vSHARP Adaptive Learned 1 1 0.881 $\pm$ 0.049 31.92 $\pm$ 2.91 0.035 $\pm$ 0.019 0.873 $\pm$ 0.051 31.73 $\pm$ 2.79 0.036 $\pm$ 0.018 0.861 $\pm$ 0.054 31.35 $\pm$ 2.75 0.039 $\pm$ 0.020
vSHARP Equi NA Learned 1 1 0.879 $\pm$ 0.050 31.62 $\pm$ 2.98 0.038 $\pm$ 0.021 0.868 $\pm$ 0.053 31.51 $\pm$ 2.88 0.038 $\pm$ 0.020 0.848 $\pm$ 0.057 30.93 $\pm$ 2.72 0.043 $\pm$ 0.021
vSHARP Equi NA Learned 1 1 0.881 $\pm$ 0.050 31.75 $\pm$ 3.00 0.037 $\pm$ 0.020 0.869 $\pm$ 0.054 31.62 $\pm$ 2.91 0.037 $\pm$ 0.020 0.849 $\pm$ 0.057 31.05 $\pm$ 2.74 0.042 $\pm$ 0.020
vSHARP Optimized Learned 1 1 0.886 $\pm$ 0.047 32.17 $\pm$ 2.87 0.033 $\pm$ 0.018 0.879 $\pm$ 0.048 31.96 $\pm$ 2.80 0.034 $\pm$ 0.018 0.870 $\pm$ 0.050 31.71 $\pm$ 2.72 0.036 $\pm$ 0.018
VarNet Adaptive Learned 1 1 0.837 $\pm$ 0.056 30.92 $\pm$ 2.40 0.041 $\pm$ 0.015 0.832 $\pm$ 0.057 30.77 $\pm$ 2.40 0.043 $\pm$ 0.016 0.829 $\pm$ 0.058 30.66 $\pm$ 2.38 0.044 $\pm$ 0.016
vSHARP Adaptive OFILK NA 1 NA 0.868 $\pm$ 0.049 31.53 $\pm$ 2.68 0.037 $\pm$ 0.019 0.856 $\pm$ 0.052 31.15 $\pm$ 2.65 0.041 $\pm$ 0.019 0.842 $\pm$ 0.056 30.70 $\pm$ 2.62 0.045 $\pm$ 0.021
vSHARP Adaptive OFTVL1 NA 1 NA 0.873 $\pm$ 0.045 30.01 $\pm$ 2.78 0.055 $\pm$ 0.031 0.861 $\pm$ 0.049 29.55 $\pm$ 2.76 0.061 $\pm$ 0.034 0.848 $\pm$ 0.053 29.01 $\pm$ 2.77 0.069 $\pm$ 0.038
vSHARP Adaptive DEMONS NA 1 NA 0.873 $\pm$ 0.050 30.36 $\pm$ 2.80 0.049 $\pm$ 0.024 0.863 $\pm$ 0.053 30.19 $\pm$ 2.78 0.051 $\pm$ 0.025 0.852 $\pm$ 0.056 29.93 $\pm$ 2.73 0.054 $\pm$ 0.026
vSHARP Equi NA OFILK NA 1 NA 0.877 $\pm$ 0.046 31.82 $\pm$ 2.74 0.035 $\pm$ 0.018 0.861 $\pm$ 0.050 31.43 $\pm$ 2.65 0.038 $\pm$ 0.018 0.838 $\pm$ 0.055 30.72 $\pm$ 2.50 0.044 $\pm$ 0.019
vSHARP Equi NA OFTVL1 NA 1 NA 0.882 $\pm$ 0.044 30.40 $\pm$ 2.99 0.051 $\pm$ 0.031 0.868 $\pm$ 0.047 30.06 $\pm$ 2.87 0.054 $\pm$ 0.031 0.847 $\pm$ 0.052 29.32 $\pm$ 2.72 0.064 $\pm$ 0.035
vSHARP Equi NA DEMONS NA 1 NA 0.880 $\pm$ 0.048 30.55 $\pm$ 2.83 0.047 $\pm$ 0.024 0.865 $\pm$ 0.052 30.32 $\pm$ 2.77 0.050 $\pm$ 0.024 0.845 $\pm$ 0.056 29.89 $\pm$ 2.64 0.054 $\pm$ 0.025
@@
Table 2: Quantitative results for various configurations on the cardiac cine test set (under unified sampling settings).
@@
Reconstruction Model Sampling Type Sampling Initialization Registration Module Joint Loss Loss Weights 4×4$\times$4 × 6×6$\times$6 × 8×8$\times$8 ×

### v2

@@ C.2 Additional Tables
$\mathcal{L}_{\mathrm{sim}},\mathcal{L}_{\mathrm{reg}}$ $\alpha$ $\beta$ $\mathcal{L}_{\mathrm{smooth}}$ E2E Loss SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$)
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ $\checkmark$ $0.960\pm 0.016$ $38.70\pm 2.43$ $0.009\pm 0.004$ $0.949\pm 0.020$ $37.52\pm 2.46$ $0.011\pm 0.006$ $0.939\pm 0.024$ $36.57\pm 2.47$ $0.014\pm 0.007$ 10.16
Voxelmorph vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ $\checkmark$ $0.948\pm 0.020$ $37.65\pm 2.48$ $0.011\pm 0.006$ $0.934\pm 0.026$ $36.28\pm 2.50$ $0.015\pm 0.008$ $0.920\pm 0.030$ $35.16\pm 2.53$ $0.020\pm 0.011$ 10.74
Transfmorph vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ $\checkmark$ $0.953\pm 0.019$ $37.81\pm 2.55$ $0.010\pm 0.005$ $0.942\pm 0.023$ $36.79\pm 2.59$ $0.013\pm 0.007$ $0.931\pm 0.028$ $35.90\pm 2.62$ $0.016\pm 0.008$ 10.11
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 3.0 $\checkmark$ $\checkmark$ $0.900\pm 0.034$ $31.48\pm 2.47$ $0.045\pm 0.020$ $0.894\pm 0.036$ $31.35\pm 2.47$ $0.046\pm 0.021$ $0.888\pm 0.039$ $31.21\pm 2.47$ $0.048\pm 0.022$ 10.90
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 2.0 $\checkmark$ $\checkmark$ $0.923\pm 0.029$ $33.67\pm 2.69$ $0.028\pm 0.014$ $0.914\pm 0.033$ $33.33\pm 2.68$ $0.030\pm 0.015$ $0.906\pm 0.035$ $33.00\pm 2.69$ $0.032\pm 0.016$ 10.40
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.5 2.0 $\checkmark$ $\checkmark$ $0.950\pm 0.019$ $37.24\pm 2.55$ $0.012\pm 0.006$ $0.940\pm 0.024$ $36.34\pm 2.56$ $0.015\pm 0.008$ $0.930\pm 0.027$ $35.56\pm 2.59$ $0.018\pm 0.009$ 11.78
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 0.0 1.0 $\checkmark$ $\checkmark$ $0.479\pm 0.085$ $22.30\pm 2.32$ $0.353\pm 0.119$ $0.475\pm 0.085$ $22.28\pm 2.32$ $0.354\pm 0.117$ $0.471\pm 0.086$ $22.26\pm 2.30$ $0.355\pm 0.116$ 10.69
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 2.0 1.0 $\checkmark$ $\checkmark$ $0.971\pm 0.012$ $40.95\pm 2.45$ $0.005\pm 0.003$ $0.961\pm 0.016$ $39.38\pm 2.48$ $0.007\pm 0.004$ $0.951\pm 0.020$ $38.13\pm 2.55$ $0.010\pm 0.005$ 10.87
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ $0.970\pm 0.012$ $40.87\pm 2.44$ $0.005\pm 0.003$ $0.959\pm 0.016$ $39.11\pm 2.47$ $0.008\pm 0.004$ $0.950\pm 0.020$ $37.93\pm 2.48$ $0.010\pm 0.006$ 10.89
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $0.972\pm 0.011$ $41.26\pm 2.44$ $0.005\pm 0.002$ $0.962\pm 0.015$ $39.54\pm 2.48$ $0.007\pm 0.004$ $0.952\pm 0.019$ $38.19\pm 2.53$ $0.010\pm 0.005$ 10.99
Learned vSHARP Adaptive $\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ $\checkmark$ $0.965\pm 0.014$ $39.68\pm 2.46$ $0.007\pm 0.003$ $0.954\pm 0.018$ $38.37\pm 2.50$ $0.009\pm 0.005$ $0.944\pm 0.022$ $37.29\pm 2.52$ $0.012\pm 0.006$ 11.91
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim2D}}$ 1.0 1.0 $\checkmark$ $\checkmark$ $0.942\pm 0.022$ $36.16\pm 2.49$ $0.015\pm 0.007$ $0.932\pm 0.026$ $35.41\pm 2.49$ $0.018\pm 0.008$ $0.922\pm 0.029$ $34.75\pm 2.48$ $0.021\pm 0.010$ 10.47
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ $0.961\pm 0.015$ $39.23\pm 2.39$ $0.008\pm 0.004$ $0.950\pm 0.020$ $37.82\pm 2.47$ $0.011\pm 0.006$ $0.939\pm 0.024$ $36.70\pm 2.51$ $0.014\pm 0.008$ 10.66
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 2.0 $\checkmark$ $0.931\pm 0.027$ $35.36\pm 2.57$ $0.019\pm 0.010$ $0.922\pm 0.030$ $34.84\pm 2.58$ $0.021\pm 0.012$ $0.913\pm 0.034$ $34.35\pm 2.57$ $0.024\pm 0.013$ 10.40
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 2.0 1.0 $\checkmark$ $0.971\pm 0.012$ $41.04\pm 2.43$ $0.005\pm 0.002$ $0.961\pm 0.016$ $39.44\pm 2.49$ $0.007\pm 0.004$ $0.951\pm 0.020$ $38.15\pm 2.53$ $0.010\pm 0.005$ 10.95
Learned vSHARP Adaptive $\checkmark$ $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ $\checkmark$ $0.964\pm 0.014$ $39.59\pm 2.43$ $0.007\pm 0.003$ $0.953\pm 0.018$ $38.29\pm 2.47$ $0.009\pm 0.005$ $0.943\pm 0.023$ $37.19\pm 2.51$ $0.012\pm 0.006$ 10.18
Learned vSHARP Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ $\checkmark$ $0.970\pm 0.012$ $40.80\pm 2.41$ $0.005\pm 0.002$ $0.954\pm 0.019$ $38.52\pm 2.51$ $0.009\pm 0.004$ $0.934\pm 0.026$ $36.54\pm 2.45$ $0.014\pm 0.007$ 9.93
Learned vSHARP kt-Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ $\checkmark$ $0.971\pm 0.011$ $40.73\pm 2.40$ $0.005\pm 0.002$ $0.952\pm 0.019$ $38.14\pm 2.44$ $0.010\pm 0.005$ $0.928\pm 0.028$ $35.92\pm 2.43$ $0.016\pm 0.008$ 9.91
Learned vSHARP Optimized $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ $\checkmark$ $0.959\pm 0.017$ $38.60\pm 2.52$ $0.009\pm 0.005$ $0.947\pm 0.021$ $37.29\pm 2.57$ $0.012\pm 0.006$ $0.936\pm 0.026$ $36.28\pm 2.57$ $0.015\pm 0.008$ 10.85
Learned VarNet Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ $\checkmark$ $0.862\pm 0.051$ $31.39\pm 2.56$ $0.046\pm 0.022$ $0.834\pm 0.055$ $30.35\pm 2.47$ $0.057\pm 0.025$ $0.814\pm 0.056$ $29.65\pm 2.32$ $0.066\pm 0.026$ 10.59
OptFlowILK vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ NA $0.971\pm 0.011$ $41.05\pm 2.42$ $0.005\pm 0.002$ $0.961\pm 0.016$ $39.35\pm 2.46$ $0.007\pm 0.004$ $0.951\pm 0.020$ $38.09\pm 2.49$ $0.010\pm 0.005$ 27.27
OptFlowTVL1 vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ NA $0.971\pm 0.011$ $41.04\pm 2.43$ $0.005\pm 0.002$ $0.961\pm 0.016$ $39.34\pm 2.47$ $0.007\pm 0.004$ $0.951\pm 0.020$ $38.09\pm 2.50$ $0.010\pm 0.005$ 30.20
DEMONS vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ NA $0.971\pm 0.011$ $41.06\pm 2.42$ $0.005\pm 0.002$ $0.960\pm 0.016$ $39.33\pm 2.47$ $0.007\pm 0.004$ $0.951\pm 0.019$ $38.10\pm 2.49$ $0.010\pm 0.005$ 29.31
OptFlowILK vSHARP Adaptive $\checkmark$ $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ NA $0.974\pm 0.010$ $41.73\pm 2.41$ $0.004\pm 0.002$ $0.963\pm 0.015$ $39.90\pm 2.48$ $0.006\pm 0.003$ $0.953\pm 0.019$ $38.50\pm 2.49$ $0.009\pm 0.005$ 27.18
OptFlowTVL1 vSHARP Adaptive $\checkmark$ $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ NA $0.974\pm 0.010$ $41.73\pm 2.43$ $0.004\pm 0.002$ $0.963\pm 0.015$ $39.90\pm 2.47$ $0.006\pm 0.003$ $0.952\pm 0.019$ $38.50\pm 2.49$ $0.009\pm 0.005$ 30.16
DEMONS vSHARP Adaptive $\checkmark$ $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ NA $0.974\pm 0.010$ $41.76\pm 2.43$ $0.004\pm 0.002$ $0.963\pm 0.015$ $39.89\pm 2.48$ $0.006\pm 0.003$ $0.952\pm 0.019$ $38.50\pm 2.49$ $0.009\pm 0.005$ 28.92
OptFlowILK vSHARP Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ NA $0.978\pm 0.009$ $42.78\pm 2.45$ $0.003\pm 0.001$ $0.962\pm 0.016$ $39.84\pm 2.51$ $0.006\pm 0.003$ $0.942\pm 0.023$ $37.41\pm 2.46$ $0.011\pm 0.006$ 26.58
OptFlowTVL1 vSHARP Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ NA $0.978\pm 0.009$ $42.78\pm 2.45$ $0.003\pm 0.001$ $0.962\pm 0.016$ $39.84\pm 2.51$ $0.006\pm 0.003$ $0.942\pm 0.023$ $37.41\pm 2.46$ $0.011\pm 0.006$ 30.08
DEMONS vSHARP Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ NA $0.978\pm 0.009$ $42.78\pm 2.45$ $0.003\pm 0.001$ $0.962\pm 0.016$ $39.84\pm 2.51$ $0.006\pm 0.003$ $0.942\pm 0.023$ $37.41\pm 2.46$ $0.011\pm 0.006$ 29.00
OptFlowILK vSHARP kt-Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ NA $0.979\pm 0.008$ $42.92\pm 2.42$ $0.003\pm 0.001$ $0.960\pm 0.015$ $39.53\pm 2.45$ $0.007\pm 0.003$ $0.938\pm 0.024$ $36.91\pm 2.40$ $0.013\pm 0.006$ 26.46
OptFlowTVL1 vSHARP kt-Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ NA $0.979\pm 0.008$ $42.92\pm 2.42$ $0.003\pm 0.001$ $0.960\pm 0.015$ $39.53\pm 2.45$ $0.007\pm 0.003$ $0.938\pm 0.024$ $36.91\pm 2.40$ $0.013\pm 0.006$ 29.74
DEMONS vSHARP kt-Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $\checkmark$ NA $0.979\pm 0.008$ $42.92\pm 2.42$ $0.003\pm 0.001$ $0.960\pm 0.015$ $39.53\pm 2.45$ $0.007\pm 0.003$ $0.938\pm 0.024$ $36.91\pm 2.40$ $0.013\pm 0.006$ 28.78
@@
Table C2: Reconstruction quantitative results for various configurations on the cardiac cine test set (under phase-specific settings).
@@
Registration Model Reconstruction Model Sampling Type Sampling Initialization Loss Details 4×4$\times$4 × 6×6$\times$6 × 8×8$\times$8 × Inference Time